The Vaughn Company issued $360,000 of 7% bonds on January 1, 2017. The bonds are due January 1, 2022, with interest payable each July 1 and January 1. The bonds were issued at 104.

Prepare the journal entries for (a) January 1, (b) July 1, and (c) December 31. Assume The Vaughn Company records straight-line amortization semiannually.
Answer #1

Premium on bonds = ($360,000 * 104%) - ($360,000)

= $374,400 - $360,000 = $14,400

Period = 5 years

Total payments = 5 * 2 = 10 payments

Amortized payment of premium per interest period = $14,400 / 10 = $1,440

Journal entries

January 1, 2017,

Cash account...........................................Debit $374,400

To Premium on bonds payable account..Credit $14,400

To Bonds payable account.......................Credit $360,000

July 1st, 2017

Interest Expense (360,000*7%) a/c...........Debit $11,160

Premium on bonds payable a/c ................Debit $1,440

To Cash account.......................................Credit $12,600

January 1st 2018,

Interest Expense (360,000*7%) a/c...........Debit $11,160

Premium on bonds payable a/c ................Debit $1,440

To Cash account.......................................Credit $12,600